OpenStudy (anonymous):

4a - b = a + 1 Solve the following equation for a.

OpenStudy (anonymous):

you have to get a in one side.

OpenStudy (anonymous):

4a - b = a + 1 4a-b-1=a subtract 1 From both sides

OpenStudy (anonymous):

Is that it when you solve for a

OpenStudy (anonymous):

yes, tim's right.....again :P

OpenStudy (anonymous):

ok i got it now 4a - b = a + 1 -b=3a+1 sub 4a to both sides -b-1=3a sub 1 TBS a=-b/3-1/3 divide both sides by 3 :D